OpenStudy (anonymous):

no matter how many times i try figuring this problem out i always mes up somewhere!! :( solve for p: –6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (saifoo.khan):

\[ –6p + 7 = 3(2p – 3) – 4(–10 + 4p)\] \[-6p+7 = 6p-9+40-16p\] \[-6p+7=-10p+31\] Shuffle. \[-6p+10p=31-7\] \[4p = 24\]\[p = 6\]

OpenStudy (saifoo.khan):

Thanks! ^_^

OpenStudy (anonymous):

Thank you!!

OpenStudy (saifoo.khan):

No problem!!